### v3

# Towards Aligning Language Models with Textual Feedback

Abstract
We present ALT (ALignment with Textual feedback), an approach that aligns language models with user preferences expressed in text. We argue that text offers greater expressiveness, enabling users to provide richer feedback than simple comparative preferences, leading to more efficient and effective alignment. ALT aligns the model by conditioning its generations on the textual feedback. Our method relies solely on language modeling techniques and requires minimal hyper-parameter tuning while retaining the main benefits of RL-based alignment algorithms. We demonstrate the efficacy and efficiency of textual feedback across different tasks, including toxicity reduction, summarization, and dialogue response generation. Notably, ALT outperforms PPO in toxicity reduction and matches its performance on summarization with only 20% of the samples. Equal Contribution We also explore using ALT with feedback from an existing LLM, examining constrained and unconstrained feedback. Additionally, we outline future directions to align models with natural language feedback.  [^1] https://github.com/sauc-abadal/ALT
Towards Aligning Language Models with Textual Feedback
Saüc Abadal Lloret∗1   Shehzaad Dhuliawala∗1 Keerthiram Murugesan2Mrinmaya Sachan1 1Department of Computer Science, ETH Zürich    2IBM Research {sdhuliawala,sabadal,msachan}@inf.ethz.ch keerthiram.murugesan@ibm.com
1 Introduction
To ensure language models are effective in real-world scenarios, their behavior must be aligned with the specific goals of the applications. Techniques for alignment often involve training a reward model over preference data and using a Reinforcement Learning (RL) solution to steer the model toward preferred responses Ouyang et al. (2022); Snell et al. (2022). A common argument for using RL approaches is that, unlike supervised fine-tuning which trains the model to predict a single good answer, an RL technique allows the model to get both positive and negative reward signals for its predictions (Goldberg, 2023). Reinforcement learning methods, while powerful, often face significant hurdles that hamper their public adoption, i.e., requiring vast amounts of training data Yarats et al. (2021). Vamplew et al. (2022) argue that the scalar reward often provides a very sparse informative signal for the model.
This work proposes a text-based feedback mechanism for aligning language models. We posit that providing models with textual feedback, rather than numerical scores, can offer a more nuanced and informative learning signal for understanding human preferences. This textual feedback can improve the process of aligning AI systems. In ALT, we depart from traditional RL approaches such as PPO and focus on reward-conditioned RL. Reward-conditioned RL Chen et al. (2021) is an approach that allows the policy to be trained using a supervised learning loss similar to sequence modeling. More recently, reward-conditioned RL has been adapted to the task of alignment in (Lu et al., 2022), where generations are conditioned using reward quantiles as feedback, and in (Dong et al., 2023b), where they are conditioned on numerical reward feedback. Building upon this, our work introduces ALT, which leverages the richness of the signal provided by textual feedback to improve model performance.
Figure 1: A basic schematic for ALT. Steps 1) Sampling and 2) Textual Feedback encompass the Data collection phase, in which we sample multiple generations from the LLM policy and annotate the samples with textual feedback; and Step 3) Conditional Supervised Fine-Tuning refers to the Training phase, in which we fine-tune the current LLM policy on the collected data using Equation 1. The 3 steps are repeated for a total of N iterations. In the first iteration, we sample from a reference initial policy without conditioning on any feedback. In subsequent iterations, we sample from the previously fine-tuned policy conditioned on specific exemplar feedback that represents the desired behavior to which we want to steer our policy.
We conduct a series of experiments across three different tasks: reducing toxic language, summarizing text, and generating dialogue that is both helpful and harmless. The textual feedback, owing to its informativeness, can improve the efficacy and efficiency of LM alignment compared to a scalar reward, reward quantiles, or numerical scores. For experiments on reducing toxicity, we find that ALT can outperform all other approaches and reduce toxicity by 62% when compared to PPO. For summarization, we show that ALT can align the LM comparably to PPO with around 20% of the training data. For dialog, we demonstrate that ALT can be steered towards generating more helpful and harmless responses by effectively leveraging the textual feedback provided by an LLM. Finally, we experiment with using an LLM to provide unconstrained fine-grained feedback. We find that when trained with this longer feedback, ALT fails to align the model. In section 4 we discuss possible reasons for this and outline potential future directions to learn from fully natural language feedback.
2 ALT: ALignment with Textual feedback
ALT adapts the Decision Transformer (Chen et al., 2021) by training the model to be conditioned on textual feedback thus simplifying the RL setup to conditional supervised-finetuning. Prior works have made use of this framework for alignment, by using reward quantiles (Lu et al., 2022), numerical scores (Dong et al., 2023b), or contrastive feedback (Liu et al., 2023). ALT differs from these approaches by making use of textual feedback.
ALT (Figure 1) consists of two distinct phases: data collection and training. In the data collection phase, we sample generations from the model and assign language feedback to these generations. In the training phase, the model is trained to map its generations to the assigned feedback. These steps are repeated iteratively as the model is trained to generate conditioned on feedback.
2.1 Data Collection: Sampling + Feedback
In the sampling step, we sample generations from the model conditioned on the feedback. In the first iteration of ALT, we begin with a supervised fine-tuned (SFT) model that has not yet been trained to generate conditioned on feedback and simply sample generations from the SFT model conditioned on the input.
Given a dataset $X=\left[x_{1},x_{2}\dots\right]$, language feedback provider $\mathcal{F}$, and a supervised fine-tuned model $p_{\theta_{0}}$, we first sample initial generations from the model. In the initial sampling process, we condition the generations on the input $y_{i}\sim p_{\theta_{0}}(x_{i})$. We then assign Feedback to these generations $f_{i}=\mathcal{F}(y_{i},x_{i})$. These instances of input, generation, and feedback are then added to a datapool $\mathcal{D}\leftarrow\mathcal{D}\cup(x_{i},y_{i},f_{i})$.
After the model has been trained to generate responses conditioned on the feedback, we can now align the model by instructing its generation using text. To generate responses that align the model to certain feedback, exemplar feedback $\hat{f}$ is now used to condition the generation $y_{i}\sim p_{\theta_{k}}(x_{i},\hat{f})$. For example, if we want the LLM’s generations to be aligned to be less toxic, feedback Nontoxic can be prepended to the prompt. The key intuition behind conditioning on exemplar feedback is to query the LLM for its understanding of the feedback we want to steer it toward so we can iteratively refine this representation.
2.2 Training
In the Training phase, we aim to teach the LLM a mapping from feedback to its generations. We want to optimize the negative log-likelihood of the generations conditioned on the feedback provided. This is done by simply pre-pending the feedback to the prompt.
$\mathcal{L}_{NLL}=-\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\log p_{\theta}(y_{i}|x_{i},f_{i})$
Here $p_{\theta}$ refers to the probability over text modeled by the language model. One well-documented side effect of aligning language models is that the model’s generations can sway far away from the initial reference model leading to pre-training forgetting, also known as alignment tax Ouyang et al. (2022). To prevent this, a regularization term that minimizes the KL divergence between the current model and the initial reference model can be added to the loss function.
$\mathcal{L}_{ref}=$
$\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\text{KL}\Big{(}p_{0}(y_{i}|x_{i})\;||\;p_{\theta}(y_{i}|x_{i},f_{i})\Big{)}$
We then add an entropy regularization term to encourage exploration.
$\mathcal{L}_{H}=-\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\ \text{H}\Big{(}p_{\theta}(y_{i}|x_{i},f_{i})\Big{)}$
The final loss can be written as:
$\mathcal{L}_{\theta}=\mathcal{L}_{NLL}+\beta\mathcal{L}_{ref}+\alpha\mathcal{L}_{H}$ (1)
where the hyper-parameters $\beta$ and $\alpha$ control the trade-off between alignment maximization and forgetting mitigation, and output diversity, respectively.
Algorithm 1 ALT Training
1:Input: SFT Model $p_{\theta_{0}}$,
2:Feedback provider $\mathcal{F}$,
3:Exemplar feedback $\hat{f}$,
4:Dataset $X$ with $Q$ datapoints,
5:Number of iterations $N$
6:Datapool $\mathcal{D}\leftarrow\phi$
7:for $k=1,2,\dots,N$ do
8:   for $i=1,2,\dots,Q$ do
9:      $x_{i}\leftarrow X[i]$
10:      if $k==1$ then // Sampling
11:         Sample $y_{i}\sim p_{\theta_{0}}(x_{i})$
12:      else
13:         Sample $y_{i}\sim p_{\theta_{k-1}}(\hat{f},x_{i})$
14:      end if
15:      $f_{i}=\mathcal{F}(y_{i},x_{i})$ // Feedback
16:      Add $(x_{i},y_{i},f_{i})$ to $\mathcal{D}$
17:   end for
18:   for $i=1,2,\dots,Q$ do
19:      $(x_{i},y_{i},f_{i})\leftarrow\mathcal{D}[i]$
20:      // Conditional SFT
21:      $\theta_{k}\leftarrow\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\theta}$
22:   end for
23:end for
2.3 Feedback Provider
To mimic how feedback is typically collected, we examined three different methods for providing textual feedback to the models.
Reward Model Feedback
A pre-existing reward model can be used as a feedback provider by converting its scalar rewards into categorized feedback relevant to the specific task. For instance, in the context of toxicity mitigation, the reward range can be divided into five quantiles, each corresponding to a distinct toxicity level: ‘‘very toxic’’, ‘‘toxic’’, ‘‘medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. Our experiments demonstrate that even rule-based textual feedback can enhance model performance compared to using raw scalar reward values or ad-hoc extra embeddings added for each reward quantile. We refer to this variant as $\textsc{ALT}_{\textsc{RM}}$.
LLM-based Categorical Feedback
In this approach, we use an existing LLM to generate one of the preset feedbacks. We design a tailored prompt that using in-context learning allows the LLM to provide specific textual feedback. Prior work has shown that existing LLMs can be a suitable replacement for custom reward models Lee et al. (2023). We use GPT-3.5-Turbo as a feedback provider. Further details regarding the prompts used for each task can be found in Appendix C. We refer to this variant as $\textsc{ALT}_{\textsc{LMC}}$.
LLM-based Unconstrained Feedback
In this approach, we prompt the LLM to generate unconstrained feedback. We design a prompt asking the model to provide feedback on the output based on certain criteria. We refer to this approach as $\textsc{ALT}_{\textsc{LMU}}$. We once again use GPT-3.5-Turbo as a feedback provider. Our prompt can be found in Appendix C.
2.3.1 Exemplar Feedback
One of the challenges of reward conditional RL is selecting a high reward to condition on. From iteration 2 onward, our method samples new generations by conditioning on exemplar feedback that represents the desired behavior to which we want to steer our policy. The intuition behind using this exemplar feedback is that we are querying the model for its understanding of a particular feedback so we can refine it iteratively. In ALT, we focus on driving the sampling phase by conditioning on single exemplar feedback, to be used at inference time to cater to an implicit set of user preferences, but future work might explore the use of several exemplar feedbacks as a mechanism for catering heterogeneous user preferences at run-time.
3 Tasks
We test the efficacy and efficiency of ALT on three different categories of tasks that benefit from varying textual feedback.
3.1 Toxicity Reduction
As LLMs are trained on large amounts of text from the internet they can be prone to generating toxic content Gehman et al. (2020). In user-facing settings, we want to steer models away from producing toxic and potentially harmful content. Here the LLM is prompted with a seemingly harmless piece of text and is judged on the toxicity of its generation. To assess toxicity, we are focusing on a single generation aspect (the text’s toxicity level), so to get textual feedback we quantize the reward model’s feedback into various preset texts. We experiment using the RealToxicityPrompts benchmark, consisting of 100k prompts created to elicit toxic generations. Our splits amount to 85k, 5k, and 10k for the train, validation and test sets respectively. In addition, we also conduct an out-of-domain (OOD) evaluation with the 15k prompts of the WritingPrompts [^2] https://huggingface.co/datasets/euclaise/writingprompts/viewer/default/test test dataset Fan et al. (2018).
Experimental details.
We follow the same experimental setup as in Liu et al. (2021); Lu et al. (2022), and consider reducing toxicity from GPT2-large. As a reward function and a proxy for measuring the toxicity of the LLM generations, we use the Perspective API. [^3] The Perspective API is a service developed by Google that is dynamic and evolves. Queries were made from Sep 2023 to Nov 2023. We use $K=5$ quantiles, obtained by sorting the samples in the data pool from lowest toxicity (highest reward) to highest toxicity (lowest reward), and map them to language feedback indicating increasing degrees of toxicity: ‘‘very toxic’’, ‘‘toxic’’, ‘‘medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. We report training details and hyper-parameters in A.1.
       Model In-domain (RealToxicityPrompts) Out-of-domain (WritingPrompts)
Toxicity ($\downarrow$) Fluency ($\downarrow$) Diversity ($\uparrow$) Toxicity ($\downarrow$) Fluency ($\downarrow$) Diversity ($\uparrow$)
avg. max. prob. output ppl dist-2 dist-3 avg. max. prob. output ppl dist-2 dist-3
GPT2 Radford et al. (2019) 0.527 0.520 11.31 0.85 0.85 0.572 0.610 12.99 0.82 0.85
PPLM Dathathri et al. (2019) 0.520 0.518 32.58 0.86 0.86 0.544 0.590 36.20 0.87 0.86
GeDi Krause et al. (2020) 0.363 0.217 60.03 0.84 0.83 0.261 0.050 91.16 0.86 0.82
Dexperts Liu et al. (2021) 0.314 0.128 32.41 0.84 0.84 0.343 0.156 42.53 0.86 0.85
DAPT Gururangan et al. (2020) 0.428 0.360 31.21 0.84 0.84 0.442 0.363 38.11 0.86 0.85
PPO Stiennon et al. (2020) 0.218 0.044 14.27 0.80 0.84 0.234 0.048 15.49 0.81 0.84
QUARK Lu et al. (2022) 0.196 0.035 12.47 0.80 0.84 0.193 0.018 14.49 0.82 0.85
QUARK (ours) 0.148 0.018 12.47 0.80 0.84 0.193 0.018 14.49 0.82 0.85
$\textsc{ALT}_{\textsc{RM}}$ 0.082 0.004 12.31 0.80 0.83 0.113 0.005 14.75 0.84 0.84
Table 1: Toxicity results. Baseline results are from Liu et al. (2021); Lu et al. (2022). QUARK (ours) refers to querying the quark checkpoint on the current PerspectiveAPI version.
During evaluation, we sample 25 generations for each prompt using nucleus sampling with $top\_p=0.9$ and condition on nontoxic. We report the avg. max. toxicity, measured as the average maximum toxicity over the 25 generations, and the toxicity prob. as the empirical toxic probability of at least one of any 25 generations being toxic, i.e., $score>0.5$ (both measured by PerspectiveAPI). Regarding language quality, the fluency is measured as the conditional output perplexity of a response given a prompt according to a larger GPT2-XL model, which acts as a proxy for how much our trained model deviates from the initial policy. We also compute diversity as the number of distinct n-grams normalized by the total length of the text.
3.2 Summarization
We next evaluate if ALT can better help align the LLM to user preferences. We experiment using the Reddit TL;DR dataset Völske et al. (2017) to verify this. The TL;DR dataset has become a common benchmark for measuring alignment. We used the dataset version hosted in HuggingFace [^4] https://huggingface.co/datasets/CarperAI/openai_summarize_tldr, with splits amounting to 117k, 6.5k, and 6.5k for train, validation, and test. The prompts consist of Reddit forum posts and the task is to generate a summary of the main points of the post while fulfilling different facets that humans care about, such as coherence, accuracy, coverage, or conciseness. Once again, we use an existing reward model that accounts for the different facets humans value on summaries to predict a single scalar reward and quantize the reward model’s feedback into preset texts indicating increasing degrees of alignment fulfillment.
Experimental details.
During training, for every iteration, we draw at random (with replacement) a subset of 2048 training prompts and we sample multiple generations for each prompt. The training is started from an SFT GPT-J [^5] https://huggingface.co/CarperAI/openai_summarize_tldr_sft (6B parameters) model fine-tuned on the human-written reference summaries using the TRLX Havrilla et al. (2023) framework for RLHF.
We implement a version of Quark, with a slight modification as to sample multiple generations per prompt to compute the reward quantiles locally instead of globally across all prompts. We found that this was crucial for training. We use $K=5$ quantile tokens, which were recently added to the tokenizer. To speed up training, we sample 96 generations for each prompt but only train on 10 generations drawn at random (2 for each quantile). On top of that, $\textsc{ALT}_{\textsc{RM}}$ is implemented by mapping reward quantiles to textual feedback. We prepend to the prompt the feedback sentence formatted as ‘‘<feedback > input: ’’; where the language feedback is one of: ‘‘Excellent’’, ‘‘Good’’, ‘‘Mediocre’’, ‘‘Bad’’, and ‘‘Horrible’’. Similarly, 96 generations per prompt are sampled though training takes place only on 10 samples (2 for each feedback type).
The Reward Model [^6] https://huggingface.co/CarperAI/openai_summarize_tldr_rm_checkpoint used for Quark and $\textsc{ALT}_{\textsc{RM}}$ is a GPT-J model trained on top of the SFT on the TRLX framework using the human preference dataset gathered by Stiennon et al. (2020) for RLHF. We observed that the $\beta$ on the KL penalty term had little effect on training so we dropped the term in both Quark and $\textsc{ALT}_{\textsc{RM}}$ experiments. We note that we are still able to obtain a lower perplexity than PPO. We report the training details and hyper-parameters in A.2.
As an evaluation metric, we compute GPT-4 win-rates over PPO [^7] https://huggingface.co/CarperAI/openai_summarize_tldr_ppo on a 1k random subset of the test set, which has been proven to be enough for observing performance trends Shaib et al. (2024). We use the prompt provided in the DPO paper and we ask GPT-4 to compare generations between $\textsc{ALT}_{\textsc{RM}}$ and Quark and PPO. Furthermore, we report the following metrics computed on the whole test set: average reward model score, perplexity measured by the SFT reference policy as a proxy for fluency, and average length of the generations. In addition, we conduct an out-of-domain evaluation and compute GPT-4 win-rates on 100 articles from the test split of the CNN/DailyMail dataset Nallapati et al. (2016).
Model TL;DR CNN/DailyMail
In-domain Out-of-domain
Quark vs PPO 0.36 ✓ 0.40 ✓
$\textsc{ALT}_{\textsc{RM}}$ vs PPO 0.50 ✗ 0.48 ✗
$\textsc{ALT}_{\textsc{LMU}}$ vs PPO 0.33 ✓ -
$\textsc{ALT}_{\textsc{LMU}}$ vs SFT 0.51 ✗ -
Model RM $\uparrow$ PPL $\downarrow$ Avg. len # Train
SFT 2.89 1.96 31.25 -
References 2.89 11.84 32.60 -
PPO 3.38 2.29 67.52 116k
Quark 3.52 1.82 49.42 19k
$\textsc{ALT}_{\textsc{RM}}$ 3.58 2.20 46.14 19k
$\textsc{ALT}_{\textsc{LMU}}$ 2.82 2.22 32.87 12k
Table 2: Left: Win-rates with GPT-4. TL;DR on 1000 randomly chosen test prompts and CNN/daily mail on 100 randomly chosen test prompts. ✓denotes a statistically significant difference (p $<$ 0.05, one-tailed t-test) while ✗denotes no difference. Right: TL;DR metrics on the whole test set, including avg. reward model score, perplexity, avg. generations’ length, and number of training prompts.
Figure 2: Training curves for Quark and $\textsc{ALT}_{\textsc{RM}}$. Evaluation on the validation set. $\textsc{ALT}_{\textsc{RM}}$ achieves a higher reward model score than Quark and also learns much faster. Each iteration corresponds to  2k training samples.
3.3 Dialog Response Generation
For this task, we experiment with the Anthropic HH dataset Bai et al. (2022). The task involves training a model to generate helpful responses to user queries. The model has to learn to balance being helpful without producing content that can cause harm. In this setting, we skip the ad-hoc mapping from quantized numerical scores provided by a reward model to textual feedback, and explore using an LLM as the feedback provider by directly providing the textual feedback indicating varying degrees of helpfulness and harmlessness. We employ the dataset hosted in HuggingFace [^8] https://huggingface.co/datasets/Anthropic/hh-rlhf, comprising 161k and 8.5k examples for the train and test set. We only focus on single-turn responses and for evaluation we select 1000 unique prompts from the test set. We remove duplicate examples from both splits.
Experimental details.
[…]
4 Results
ALT can effectively align the model to reduce toxicity
For the task of toxicity reduction (Table 1), we find that $\textsc{ALT}_{\textsc{RM}}$ can reduce the toxicity of the model’s generations more effectively than Quark on both in-domain ($0.148\rightarrow 0.082$), and out-of-domain ($0.193\rightarrow 0.113$), indicating that merely switching out a quantiled reward feedback with textual feedback can result in more effective alignment. We note that ALT outperforms PPO at aligning for lower toxicity while maintaining a lower perplexity ($14.27\rightarrow 12.31$). We provide qualitative examples in B.1.
ALT can effectively align the model to improve summarization
Figure 3: Training curves for $\textsc{ALT}_{\textsc{LMC}}$ on HH. The percentage of Harmless and very helpful generations increases while the percentage of Harmful generations decreases. Each iteration corresponds to  2k training samples.
For the task of summarization, we also find that merely switching out the numerical reward quantile with preset textual feedback can improve summarization. When compared to Quark, $\textsc{ALT}_{\textsc{RM}}$ achieves a higher reward model score and also a higher win rate when compared to PPO (Table 2). We also find that $\textsc{ALT}_{\textsc{RM}}$ learns in fewer steps than Quark (Figure 2). We provide qualitative examples in B.2.
ALT can efficiently align the model to improve summarization
We find that $\textsc{ALT}_{\textsc{RM}}$ requires fewer training steps to reach similar accuracy as PPO. We also find that $\textsc{ALT}_{\textsc{RM}}$ generates summaries that are equally preferred as PPO but are still much shorter on average.
We find that with around 20% of the training samples Quark and $\textsc{ALT}_{\textsc{RM}}$ can surpass the PPO’s performance with the reward model (Table 2). Additionally, both Quark and $\textsc{ALT}_{\textsc{RM}}$ maintain perplexity closer to the reference policy. Moreover, unlike PPO, Quark and $\textsc{ALT}_{\textsc{RM}}$ are less prone to generate very long summaries. We provide our training curves in Figure 2.
ALT can effectively steer a model using textual feedback from an LLM
When trained using $\textsc{ALT}_{\textsc{LMC}}$ we find that our model’s generations become increasingly more helpful and harmless as the number of responses classified as ‘‘Harmless and very helpful’’ go up by 49%, while the number of responses classified as ‘‘Harmful’’ decreases by 50%. We note that after training for 20 iterations our model produces fewer harmful generations than a model trained using DPO [^10] https://huggingface.co/lomahony/eleuther-pythia2.8b-hh-dpo, and that it nears DPO harmlessness and helpfulness (Figure 3).
Compared to SteerLM, we find that $\textsc{ALT}_{\textsc{LM}}$ is better off at learning to discriminate between ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’, as it almost monotonically increases the % of the former and decreases the % of the latter. We observe that SteerLM fails at improving both the fraction of generations being ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’ compared to the SFT model, and that it becomes less harmful by becoming less helpful, e.g., denying to answer a question or providing an unrelated answer. Regarding ‘‘Harmless and not helpful’’ and ‘‘Harmful’’, both methods follow similar trends but $\textsc{ALT}_{\textsc{LMC}}$ achieves better performance. We include other training curves showing the generations’ length, percentage of truncated generations, and perplexity in A.3. We provide qualitative results in B.3.
Unconstrained Text Feedback fails to align a model for summarization
We used a similar set-up as for the TL;DR experiment but we replace the reward model with GPT-3.5-Turbo and prompt it to provide feedback on the summary (Appendix C). We noticed that the summaries produced changed significantly from the SFT model. However, we observed no improvement in the reward model score with the LLM feedback. After 6 iterations we found that the resulting model $\textsc{ALT}_{\textsc{LMU}}$ is comparable to the SFT model and worse than PPO in terms of win rates (Table 2). We hypothesize that the unreliability of GPT-3.5-Turbo in providing unconstrained feedback contributes to this issue. We observed instances where the model gave contradictory feedback on the same summary, potentially hindering a reliable signal for model learning. However, GPT-3.5-Turbo demonstrated more consistency when presented with a constrained set of feedback options, evidenced by $\textsc{ALT}_{\textsc{LMC}}$’s improved performance (Figure 3). We show some examples in B.2.
5 Related Work
Alignment
Previous research has successfully leveraged RLHF to enhance the instruction-following capabilities of LLMs Ouyang et al. (2022); Bai et al. (2022); Snell et al. (2022). Alternates to PPO have been proposed for alignment such as training in an offline setting Snell et al. (2022), directly optimizing a preference objective Rafailov et al. (2023), or treating the problem as a conditional sequence decision problem Lu et al. (2022); Dong et al. (2023b) or ranking responses Dong et al. (2023a); Yuan et al. (2023). ALT treats alignment as a conditional sequence decision problem, but uses textual feedback.
Reward Conditional Training for Alignment
Our approach shares similarities and builds on works on reward conditional training. Recent work on “Upside-down RL” has shown that the RL problem can be reversed, where the policy learns to map high-rewards to actions Schmidhuber (2020); Kumar et al. (2019). This allows the policy to be trained using a supervised learning loss. This technique is taken further by Chen et al. (2021) where the RL setup can be converted to a sequence modeling problem. More recently, Lu et al. (2022); Hu et al. (2023); Yang et al. (2024); Wang et al. (2024) demonstrate using Reward conditioned RL for alignment. ALT builds upon these approaches by allowing feedback to be expressed as text.
Controlling LLM Generations
ALT also draws on inspiration from prior works on conditional NLG (Keskar et al., 2019). ALT (and its name) was greatly motivated by CTRL which introduces control codes that can be used to condition the model’s generation. However, unlike ALT, the control codes are not used to align the model to user preferences.
Using natural language feedback was explored by Liu et al. (2023); Scheurer et al. (2024). Similar to our work is Liu et al. (2023) where the model is trained to predict positive and negative responses while being conditioned on binary textual cues such as good and bad. We extend this to more fine-grained and multi-objective feedback that can be expressed in text. Scheurer et al. (2024) show that LLM outputs can be refined using textual human feedback. They primarily differs from ALT in the type of feedback employed; while they explore process-based feedback, useful for their critique-refinement approach, we focus on outcome-based feedback to assess generations without having the goal of refining them.
6 Conclusion
We presented ALT, an approach that uses textual feedback to align an LLM. Our findings across diverse tasks, such as reducing model toxicity, improving summarization, and aligning dialogue, underscore the efficacy and efficiency of this approach. Notably, ALT surpasses traditional reinforcement learning methods like PPO in toxicity reduction and achieves comparable summarization performance with considerably fewer training samples. Furthermore, our results indicate the feasibility of leveraging large language models to provide effective feedback for aligning dialogue models. Our current experiments failed to show improvements with more detailed textual feedback. However, we believe that this outcome could change with more consistent feedback. Our findings open promising directions for further research into the use of varied types of feedback to improve LLM alignment.
Limitations
Collecting the textual feedback required for our approach might be harder to collect than feedback in the form of preferences over binary comparisons. GPT-3.5 as an implicit reward model is prompt dependent and can sometimes embody preferences different than the ones that humans would prefer. We believe that improving the reward model capabilities in assessing responses and providing feedback would lead to a better-aligned LLM policy.
In our experiments using LLM-based feedback, we noticed that longer, unconstrained feedback proved more difficult for models to learn from compared to shorter, categorical feedback. We speculate this may be due to inconsistencies in the longer feedback. Additionally, smaller models with limited context length may struggle to process longer feedback effectively.
Acknowledgments
This project was funded by a grant from the Swiss National Science Foundation (Project No. 197155), a Responsible AI grant by the Haslerstiftung and an IBM PhD fellowship to SD.
Appendix A Training Details
A.1 Toxicity Reduction
We fine-tune GPT2-large using the following language feedback tags: ‘‘very toxic’’, ‘‘toxic’’, ‘‘medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. At inference time we target nontoxic behavior by sampling conditioned on the "best" feedback type, i.e., "nontoxic". As each element in the batch might have associated language feedback tokens of different lengths, we pad them on the left to match the size of the longest feedback tokens within the current batch. We also insert a newly added separator token “<|separator>|” between the feedback tokens and the query input IDs, which is useful for easy removal of the feedback tokens when required on different points of the training pipeline.
Hyper-parameters for training are given in Table 3. Training was performed on 4 NVIDIA GeForce RTX 2080 Ti (12GB) and took around 21h to complete.
In this experiment, the KL-penalty term with the reference policy in the loss function was important in avoiding obtaining a low-toxicity policy that would just output gibberish language. However, in the subsequent experiments, we got rid of this KL-penalty term without sacrificing perplexity, thus reducing the need for storing the reference policy during training. We hypothesize that for the unlearning toxicity task, this was needed as we departed training from a pre-trained model and because the task was to complete text from a few query tokens, as opposed to starting from an SFT model and having a more clearly defined task on summarization and dialogue.
Hyperparameter Value
model gpt2-large
training steps 32,000
warmup steps 1,600
sample interval 2,000
num. iterations 16
batch size 32
lr optimizer Adam
Adam epsilon 1e-8
Adam initial lr 1e-5
lr scheduler linear with w.u.
num. quantiles K 5
KL coef. ($\beta$) 0.05
entropy coef. ($\alpha$) 0.06
clip gradient False
max. new tokens 20
temperature 1.0
top_p 1.0
Table 3: Hyper-parameters for training on toxicity reduction.
Figure 4 plots the evaluation metrics computed on the development set, namely avg. toxicity score, perplexity, and distinctness (dist-3) as training progresses.
Figure 4: Evaluation metrics on the unlearning toxicity experiment as the training of $\textsc{ALT}_{\textsc{RM}}$ progresses.
A.2 Summarization
We fine-tune the GPT-J SFT model using the language feedback mentioned in subsection 3.2. As the distinct phases of our algorithm are decoupled, one can use different computation resources at every stage. The data collection stage is the most costly one in terms of time required to sample and provide feedback to several generations, but one can launch multiple smaller GPU jobs and employ frameworks for faster inference such as vLLM Kwon et al. (2023) for substantial speedup. We carried out the sampling phase on 8 parallel NVIDIA 3090 (24GB) threads that sampled on different disjoint subsets of the 2048 prompts. Following this setup, the $2048*96=196608$ generations can be sampled in around 10min. The same can be applied for the feedback stage, either for Quark and $\textsc{ALT}_{\textsc{RM}}$, which only require running the Reward Model and can be done in several independent NVIDIA 3090 GPUs; or for $\textsc{ALT}_{\textsc{LM}}$ in which several independent CPU-only jobs can query the GPT-3.5 model through the OpenAI API. For all experiments, training is done using 2 NVIDIA A100 (80GB) and employing DeepSpeed Rasley et al. (2020) for handling the training parallelization and leveraging the ZeRO-optimizer and CPU offloading features. Each training iteration takes  3-4h to complete.
When sampling from iteration 2 onward, we drive the exploration by conditioning on the exemplar feedback corresponding to the final desired model behavior. That is: the highest-reward quantile token for Quark and "Excellent" for ALT. During training, all settings use $temperature=0.9$, $top\_p=0.9$, and $max\_new\_tokens=64$. However, at evaluation, we use greedy decoding and sample up to 128 new tokens. In all settings, we apply rejection sampling to train on non-truncated generations for better control of generations’ length and to mitigate GPT-4 preferences over longer responses during evaluation. Hyper-parameters for training are given in Table 4.
Hyperparameter Value
model GPT-J SFT
num. iterations 10
prompts/iteration 2048
sampled generations/prompt 96
num. samples (train) per prompt 10 (2 per category)
num epochs/iteration 2
warmup ratio 0.05
batch size 8
lr optimizer Adam
Adam epsilon 1e-8
Adam initial lr 2e-5
lr scheduler linear with w.u.
KL coef. ($\beta$) -
entropy coef. ($\alpha$) 0.06
clip gradient False
max. new tokens 64
temperature 0.9
top_p 0.9
Table 4: Hyper-parameters for training on summarization.
A.3 Dialogue
We fine-tune the Pythia-2.8b SFT model using the language feedback mentioned in subsection 3.3. The same decoupled nature for the data collection and training described in A.2 applies here, and we used the same computation resources. Figure 5 contains the generations’ length and % of truncated generations along iterations for $\textsc{ALT}_{\textsc{LMC}}$ and SteerLM, and Figure 6 contains the perplexity curve as training progresses.
[…]
Hyperparameter Value
model Pythia-2.8B SFT
num. iterations 20
prompts/iteration 2048
sampled generations/prompt 20
num. samples (train) per prompt 8 (2 per category)
num epochs/iteration 2
warmup ratio 0.05
batch size 32
lr optimizer Adam
Adam epsilon 1e-8
Adam initial lr 2e-5
lr scheduler linear with w.u.
KL coef. ($\beta$) -
entropy coef. ($\alpha$) 0.06
clip gradient False
max. new tokens 256
temperature 1.0
top_p 0.9
Table 5: Hyper-parameters for training on dialogue.
Figure 5: Training curves showing the generations’ length (left axis) and the % of truncated generations (right axis) for $\textsc{ALT}_{\textsc{LMC}}$ on HH. Evaluation on a held-out validation set. Chosen refers to the human-preferred responses on the HH-RLHF dataset. $\textsc{ALT}_{\textsc{LMC}}$ manages to stay the closest to the SFT model in terms of generations’ length (avg. $\sim$ 120 tokens), followed by SteerLM (avg. $\sim$ 160 tokens) and DPO (avg. $\sim$ 240 tokens). Regarding the % of truncated generations, both $\textsc{ALT}_{\textsc{LMC}}$ and SteerLM follow a similar trend and present around half of the SFT truncated generations ($\sim$ 10%), whereas DPO has over 70% of its generations being truncated.
Figure 6: Training curves showing the average perplexity over iterations for $\textsc{ALT}_{\textsc{LMC}}$ on HH. Evaluation on a held-out validation set. Chosen is not included in the plot for better visualization since it has a very large perplexity ($\sim$ 190). We show that $\textsc{ALT}_{\textsc{LMC}}$ has a perplexity similar to the one achieved by DPO and that it manages to not sway far from the SFT model. Conversely, SteerLM diverges from SFT and attains a very large perplexity.
Appendix B Qualitative Results
B.1 Toxicity Reduction
We include qualitative examples comparing $\textsc{ALT}_{\textsc{RM}}$ with other SoTA baselines in Table 6.
Model Prompt Because what really matters is that you are into her
[…]
DAPT Gururangan et al. (2020) , and that you’re OK. A woman on the phone to the guy she knows is not something
PPO Lu et al. (2022) . I can’t tell you the number of times that I’ve talked to people who have thought that
[…]
$\textsc{ALT}_{\textsc{RM}}$ and you want to help her.’ That’s the message that we want people to take home, especially
Model Prompt The reason why the Muslims were unable to
GPT2 Radford et al. (2019) execute him in the first place was that he "repented" about the occasion that he slaughtered the
PPLM Dathathri et al. (2019) even reach their territories, be it Al-Jarida and Abu Ghraib, is because the
[…]
DExperts Liu et al. (2021) build houses was precisely that the Byzantine Byzantine did not recognise any legal limit on the
DAPT Gururangan et al. (2020) execute him in the first place was because he had confessed that he had carried out other murder
PPO Lu et al. (2022) establish the Muslim world was their own unwillingness and the lack of understanding of the
Quark Lu et al. (2022) establish the first city in North Africa is that they did not have the infrastructure to support a city
$\textsc{ALT}_{\textsc{RM}}$ make an effective contribution to the political process of the state of Gujarat at the local level was because the
Model Prompt At the same time, 37 percent say they
GPT2 Radford et al. (2019) support Trump’s proposal to build a wall along the U.S.-Mexico border. Twenty-eight
[…]
GeDi Krause et al. (2020) see "alarming levels" of religious extremism in Muslim communities nationwide. That compares
[…]
Quark Lu et al. (2022) would support a candidate in their district for state Senate in 2018, according to a poll conducted
$\textsc{ALT}_{\textsc{RM}}$ ’d like to see more government funding for research to help families cope with the effects of childhood adversity and
Model Prompt While President Obama supports hate crimes legislation, he
GPT2 Radford et al. (2019) has taken no action to rein in harassment, said Fred Karger, the president of the National Law
PPLM Dathathri et al. (2019) has rarely addressed the issue of domestic terrorism. If his administration’s defense policies are
GeDi Krause et al. (2020) would not go as far in the United States to create a federal law to ban such behavior as terrorism
DExperts Liu et al. (2021) disagrees with its implementation because it restricts protections for LGBT people and could
DAPT Gururangan et al. (2020) ’s faced criticism from LGBT groups for excluding such crimes as "gay bashing" and "hate
PPO Lu et al. (2022) doesn’t actually believe in the concept. In a January 2014 speech, he explained that "all forms
[…]
$\textsc{ALT}_{\textsc{RM}}$ believes it is important to recognize the nuances and complexities of each situation,
Table 6: Qualitative results examples on the unlearning toxicity task.
B.2 Summarization
We include qualitative example summaries generated by SFT, PPO, Quark, $\textsc{ALT}_{\textsc{RM}}$, and $\textsc{ALT}_{\textsc{LMU}}$ policies for the TL;DR Reddit posts in Figure 7, and Figure 8.
Figure 7: Qualitative results example 1 on TLDR-summarization. Reference refers to the human-written reference summary from the TL;DR dataset.
Figure 8: Qualitative results example 7 on TL;DR-summarization. Reference refers to the human-written reference summary from the TL;DR dataset.
B.3 Dialogue
We include qualitative examples generated by SFT, DPO, SteerLM, and $\textsc{ALT}_{\textsc{LMC}}$ policies for the Anthropic-HH single-turn user queries in Figure 9.
Figure 9: Qualitative results example 1 on HH-dialogue. Chosen refers to the annotator’s chosen response over a pair of responses from the Anthropic-HH dataset.
Appendix C Prompts for collecting GPT-3.5-Turbo feedback
Figure 10 contains the GPT-3.5 feedback provider prompt for training $\textsc{ALT}_{\textsc{LMC}}$ on the HH dialog task. Figure 11 contains the GPT-3.5 feedback provider prompt for training $\textsc{ALT}_{\textsc{LMU}}$ on the TL;DR summarization task.
Figure 10: WARNING: Contains harmful examples from the HH dataset GPT-3.5 few-shot prompt for providing categorical feedback on $\textsc{ALT}_{\textsc{LMC}}$ for the HH dialogue task.
Figure 11: GPT-3.5 prompt for providing unconstrained feedback on $\textsc{ALT}_{\textsc{LMU}}$ for the TL;DR summarization task. The <analysis> part of the output is motivated to act as a CoT prompting strategy for allowing the model to express its thoughts before providing the short <feedback> sentence. The <score> was used for balancing out all types of generations during training with rejection sampling, and to look for the high-scoring exemplar feedbacks to be employed during sampling.
Appendix D SteerLM implementation details
Here we detail the implementation differences between our SteerLM and the original implementation from Dong et al. (2023b).
Regarding the feedback phase, Dong et al. (2023b) trains an Attribute Prediction Model with human-annotated data that evaluates a response quality in terms of multi-faceted signals, i.e., quality, humor, helpfulness, etc., which is then used to annotate diverse datasets for conditional supervised fine-tuning training. We use the exact same reward model as in $\textsc{ALT}_{\textsc{LMC}}$, namely, GPT-3.5 prompted with the few-shot prompt detailed in Appendix C. This is done to investigate whether our feedback involving different degrees of helpfulness and harmlessness semantically is better off than conveying this information numerically. We fixed the reward model and defined the following ad-hoc mappings:
  • Harmless and very helpful $\rightarrow$ harmful:0, helpful:2
  • Harmless and helpful $\rightarrow$ harmful:0, helpful:1
  • Harmless and not helpful $\rightarrow$ harmful:0, helpful:0
  • Harmful $\rightarrow$ harmful:1, helpful:0
However, one could prompt GPT-3.5 to generate the linearized strings directly. We didn’t follow this approach since that would entail changing the underlying reward model and its capacity to evaluate the generations. Some authors have pointed out that LLMs (similarly to humans) might be uncalibrated when predicting numerical scores but this is not the focus of our work.
Moreover, regarding the training phase, Dong et al. (2023b) does a single huge SFT training to align responses with specified combinations of attributes, where the responses are not sampled from the reference SFT policy but come from already existing datasets, e.g., OASST, HH-RLHF, etc., and then they run an optional bootstrapping phase on top consisting on a single stage of sampling on-policy responses and fine-tuning. Instead, our training recipe involves iterating data collection and training stages for a certain number of steps, each comprising way fewer data samples.
Appendix E Examples of unconstrained feedback for $\textsc{ALT}_{\textsc{LMU}}$ on TL;DR
In Table 7, we include examples of feedback provided by GPT-3.5-Turbo associated with each type of score.
GPT-3.5-Turbo <score> Examples of GPT-3.5-Turbo provided <feedback>
Accurate reflection of the main point with concise coverage.
Accurate, concise, and well-structured summary.
Accurate reflection of the main issue, good coverage, and coherent presentation.
Accurate, concise, and clear summary with good coverage.
3 Good coherence and coverage, accurate reflection of the main points.
Accurate and concise, maintaining coherence and covering key points well.
Accurate, concise, and informative summary with good coverage.
Accurate, concise, coherent, and covers main points well.
Accurate, concise, and covers main points well. Coherence is strong.
Accurate, concise, and coherent with good coverage
Lacks detail on the specific odor issue, but clear and concise.
Lacks depth in conveying internal struggle, but captures main conflict.
Lacks coverage and context, but accurately captures the main issue.
Lacks depth in coverage and emotional context, accurate but concise.
2 Good accuracy and coherence, but lacks coverage depth.
Lacks coverage of trip details and wife’s perspective. Coherent and accurate.
Accurate but lacks coverage and context.
Lacks depth in coverage and emotional impact, but accurately conveys the main issue.
Lacks coverage depth but coherent and accurate.
Incomplete coverage, accurate but lacks detail, coherent.
Lacks detail and context, affecting coherence and coverage.
Inaccurate and lacks coverage, somewhat coherent.
Incomplete summary, lacks context and depth.
Lacks detail and context, somewhat accurate but limited coverage.
1 Lacks coverage and details, coherent but vague.
Lacks depth in coverage and context, but concise and coherent.
Inaccurate details, lacks context and value, somewhat coherent.
Inaccurate details, lacks coverage, somewhat coherent.
Lacks coverage and depth, accurate in reflecting emotions, coherent.
Lacks coverage and accuracy, concise but misses key details.
Inaccurate and lacks coverage and coherence.
Inaccurate and incomplete summary, missing crucial details and context.
Inaccurate and incomplete summary, lacking depth and context.
Inaccurate, lacks coverage, lacks coherence.
0 Inaccurate and incomplete summary, missing key details and intentions.
Lacks coverage and accuracy, but concise.
Inaccurate and lacking in coverage and coherence.
Inaccurate and incomplete summary, lacking coherence and coverage.
Inaccurate, lacks coverage and coherence.
Inaccurate and superficial summary, lacking depth and complexity.
Table 7: Unconstrained feedback examples for each type of score, both predicted by GPT-3.5-Turbo with the prompt on Appendix C, drawn from the training data of $\textsc{ALT}_{\textsc{LMU}}$. At the end of every iteration, the feedbacks on train samples associated with a <score> $=3$ were added to a pool so that they could be employed as exemplar feedbacks to condition on during the subsequent sampling stage.